OpenStudy (anonymous):

x - 2y = 1 solve for y. Is (x - 1)/2 = y the correct answer?

OpenStudy (anonymous):

Yes, it's correct.

OpenStudy (anonymous):

x- 2y = 1 -2y = 1-x 2y = x-1 y = (x-1)/2 yes it is the correct answer

OpenStudy (anonymous):

you guys are awesome!!! Thank you so much for your help